OpenStudy (anonymous):

or this problem, the answer choices are "always", "sometimes", or "never". x • x = 2x A. Always B. Sometimes C. Never

OpenStudy (anonymous):

never

OpenStudy (anonymous):

as 1not equals 2

Directrix (directrix):

B.Sometimes x * x = 2x x^2 - 2x = 0 x(x - 2) = 0 x = 0 or x = 2 The equation given in the problem is true if x = 0 or if x = 2.

OpenStudy (anonymous):

oh yah @directrix forgot that aspect of the sum

Directrix (directrix):

And, dividing by a variable is frequently done but is dicey because if x = 0, you are dividing by 0 which is undefined. In short, you lose one of the two solutions. That's why I factor if I can and avoid dividing through by a variable.

OpenStudy (anonymous):

yup, will keep this sum in mind, tricky if you overlook it.